what interest was earned on an account that began with $42000 and paid $1000 per month for 6 years? Note: PMT & I\Y must be the same frequency= if PMT monthly then I\Y is monthly. to get annual rate =(X12)

Use BA II Plus calculator!!
Answer #1

Enter the following on the calculator

PV = -42000

PMT= 1000

N= 6*12=72

Find I/Y = 1.646%

Hence Annual rate = 1.646%*12

= 19.75%